# Marangoni-driven pattern transition and the formation of ridges and hills in surfactant-covered parametric surface waves

### v1

@@ Abstract
Our computational domain is a laterally-periodic box of area $\tilde{\lambda}_{c}\times\tilde{\lambda}_{c}$ and height $\tilde{H}$, where $\tilde{\lambda}_{c}$ is the critical wavelength. A thin film of a water-glycerine mixture (of density $\tilde{\rho}_{w}=1000~{}\rm kg/m^{3}$, viscosity $\tilde{\mu}_{w}=0.025~{}\rm kg/m~{}s$) of depth $\tilde{h}=1~{}\rm{mm}$ is considered with air as the background fluid. The interface (of surface tension $\tilde{\sigma}=70~{}\rm mN/m$) is vibrated at a frequency $f=100~{}\rm{Hz}$ and acceleration amplitude $\tilde{A}$. We assume that a layer of insoluble surfactant monomers (of concentration $\tilde{\Gamma}$) covers the interface such that the number of moles of monomers on the surface is conserved. The surfactant-dependent surface tension is parameterised by the Langmuir-Szyszkowski equation of state, $\tilde{\sigma}(\tilde{\Gamma})/\tilde{\sigma}_{0}=1+\beta_{s}\ln(1-\tilde{\Gamma}/\tilde{\Gamma}_{\infty})$, where $\tilde{\Gamma}_{\infty}$ is the maximum packing concentration and $\beta_{s}$ is the elasticity parameter reflecting the effectiveness of the surfactant in lowering $\tilde{\sigma}$.
@@
Figure 1: (a) Phase diagram in the $\beta_{s}-\Gamma_{0}$ parameter plane showing the inertia-dominated (violet) and Marangoni-dominated (pink) regions. The solid, dotted, dot-dashed, and dashed lines correspond to the $B=1,1.23,1.46,2$ contours, respectively. The four typical patterns are squares, asymmetric squares, weakly wavy stripes, and ridges and hills. The phase boundaries are accurate to within $\Delta B=\pm 0.1$. The corresponding values of $B$ and $F_{c}^{D}$ are reported in table II of the supplemental material [9]. (b) Spatiotemporal evolution of the surface deflection $\zeta$ over one time period is shown from left to right; squares ($B=0.92$), asymmetric squares ($B=1.23$), weakly wavy stripes ($B=1.51$), and ridges and hills ($B=1.83$) are shown from top to bottom rows, respectively. (c,d) $\zeta_{mn}$ and $\Gamma_{mn}$, the maximal magnitudes over time of the $\zeta$ and $\Gamma$ Fourier coefficients, respectively, as a function of $B$.
@@
We initialize our simulations with a plane interface and uniformly distributed surfactant coverage $\tilde{\Gamma}_{0}$. The interfacial dynamics are coupled to a surfactant transport equation that accounts for inertia, viscosity, gravity, capillarity, Marangoni stresses, surface diffusion, and parametric forcing [10]. Non-dimensionalization of the equations is carried out, in which $\tilde{h}$, $1/2\pi f$, and $\tilde{\Gamma}_{\infty}$ are taken as the characteristic length, time, and concentration scales, leading to six dimensionless numbers. We set these and their values to be Reynolds $Re\equiv 2\pi\tilde{\rho}_{w}\tilde{h}^{2}f/\tilde{\mu}_{w}=25.3$, Weber $We\equiv 4\pi^{2}\tilde{\rho}_{w}\tilde{h}^{3}f^{2}/\tilde{\sigma}_{0}=18.37$, Péclet, $Pe\equiv 2\pi\tilde{h}^{2}f/\mathcal{D}=2.5\times 10^{5}$, and Froude $Fr\equiv 2\pi f\sqrt{\tilde{h}/g}=6.34$ numbers. Rather than the usual Marangoni number, $Ma=\beta_{s}/We$, we introduce $B\equiv(\tilde{\sigma}_{0}-\tilde{\sigma}(\tilde{\Gamma}_{0}))/(2\pi\tilde{\mu}_{w}\tilde{h}f)$, which is the ratio of vibratory inertial and Marangoni timescales, to quantify the combined effects of $\beta_{s}$ and $\tilde{\Gamma}_{0}$; $B$ is varied systematically to assess the emergent pattern transitions at the acceleration amplitude obtained by $Fr_{t}=2\pi f\sqrt{\tilde{A}/g}$. Detailed descriptions of the formulation, scaling, and numerical method [10] are presented in Section A of the supplemental material [9].
@@
After a transient phase, Faraday waves appear, which correspond to subharmonic waves whose amplitude is steady and whose response period $T$ is twice that of the forcing period. We set $t=0$ to be an instant within the steady-amplitude Faraday-wave regime. Hereafter, we discuss our results in terms of dimensionless variables (with the tilde suppressed). The dimensionless critical wavelength $\lambda_{c}$ and acceleration amplitude $F_{c}=Fr/Fr_{t}$ for a clean interface are $5.30$ and $12.34$, respectively. We studied the effect of varying $B$ on the threshold acceleration $F_{c}^{D}$ with the details provided in Section B of the supplemental material [9]. Our results show that increasing $B$ stabilises the interface, as observed in previous studies [5, 4]. The computations for assessing the influence of $B$ on the interfacial dynamics in the nonlinear regime are then carried out for $F=1.1F_{c}^{D}$ for which square patterns are observed in the surfactant-free case.

### v2

@@ Abstract
Our computational domain is a laterally-periodic box of area $\tilde{\lambda}_{c}\times\tilde{\lambda}_{c}$ and height $\tilde{H}$, where $\tilde{\lambda}_{c}$ is the critical wavelength. A thin film of a water-glycerine mixture (of density $\tilde{\rho}_{w}=1000~{}\rm kg/m^{3}$, viscosity $\tilde{\mu}_{w}=0.025~{}\rm kg/m~{}s$) of depth $\tilde{h}=1~{}\rm{mm}$ is considered with air as the background fluid. The interface (of surface tension $\tilde{\sigma}=70~{}\rm mN/m$) is vibrated at an angular frequency $\omega=200\pi~{}\rm{rad/s}$ and acceleration amplitude $\tilde{A}$. We assume that a layer of insoluble surfactant monomers (of concentration $\tilde{\Gamma}$) covers the interface such that the number of moles of monomers on the surface is conserved. The surfactant-dependent surface tension is parameterised by the Langmuir-Szyszkowski equation of state, $\tilde{\sigma}(\tilde{\Gamma})/\tilde{\sigma}_{0}=1+\beta_{s}\ln(1-\tilde{\Gamma}/\tilde{\Gamma}_{\infty})$, where $\tilde{\Gamma}_{\infty}$ is the maximum packing concentration and $\beta_{s}$ is the elasticity parameter reflecting the effectiveness of the surfactant in lowering $\tilde{\sigma}$.
@@
Figure 1: (a) Phase diagram in the $\beta_{s}-\Gamma_{0}$ parameter plane showing the inertia-dominated (violet) and Marangoni-dominated (pink) regions. The solid, dotted, dot-dashed, and dashed lines correspond to the $B=1,1.23,1.46,2$ contours, respectively. The four typical patterns are squares, asymmetric squares, weakly wavy stripes, and ridges and hills. The phase boundaries are accurate to within $\Delta B=\pm 0.1$. The corresponding values of $B$ and $F_{c}^{N}$ are reported in table II of the supplemental material [9]. (b) Spatiotemporal evolution of the surface deflection $\zeta$ over one time period is shown from left to right; squares ($B=0.92$), asymmetric squares ($B=1.23$), weakly wavy stripes ($B=1.51$), and ridges and hills ($B=1.83$) are shown from top to bottom rows, respectively. (c,d) $\zeta_{mn}$ and $\Gamma_{mn}$, the maximal magnitudes over time of the $\zeta$ and $\Gamma$ Fourier coefficients, respectively, as a function of $B$.
@@
We initialize our simulations with a plane interface and uniformly distributed surfactant coverage $\tilde{\Gamma}_{0}$. The interfacial dynamics are coupled to a surfactant transport equation that accounts for inertia, viscosity, gravity, capillarity, Marangoni stresses, surface diffusion, and parametric forcing [10]. Non-dimensionalization of the equations is carried out, in which $\tilde{h}$, $1/\omega$, and $\tilde{\Gamma}_{\infty}$ are taken as the characteristic length, time, and concentration scales, leading to six dimensionless numbers. We set these and their values to be Reynolds $Re\equiv\tilde{\rho}_{w}\tilde{h}^{2}\omega/\tilde{\mu}_{w}=25.3$, Weber $We\equiv\tilde{\rho}_{w}\tilde{h}^{3}\omega^{2}/\tilde{\sigma}_{0}=18.37$, Péclet, $Pe\equiv\tilde{h}^{2}\omega/\mathcal{D}=2.5\times 10^{5}$, and Froude $Fr\equiv\omega\sqrt{\tilde{h}/g}=6.34$ numbers. Rather than the usual Marangoni number, $Ma=\beta_{s}/We$, we introduce $B\equiv(\tilde{\sigma}_{0}-\tilde{\sigma}(\tilde{\Gamma}_{0}))/(\tilde{\mu}_{w}\tilde{h}\omega)$, which is the ratio of vibratory inertial and Marangoni timescales, to quantify the combined effects of $\beta_{s}$ and $\tilde{\Gamma}_{0}$; $B$ is varied systematically to assess the emergent pattern transitions at the acceleration amplitude obtained by $F=\tilde{A}/g$. Detailed descriptions of the formulation, scaling, and numerical method [10] are presented in Section A of the supplemental material [9].
@@
After a transient phase, Faraday waves appear, which correspond to subharmonic waves whose amplitude is steady and whose response period $T$ is twice that of the forcing period. We set $t=0$ to be an instant within the steady-amplitude Faraday-wave regime. Hereafter, we discuss our results in terms of dimensionless variables (with the tilde suppressed). The dimensionless critical wavelength $\lambda_{c}$ and acceleration amplitude $F_{c}$ for a clean interface are $5.30$ and $12.34$, respectively. We studied the effect of varying $B$ on the threshold acceleration $F_{c}^{N}$ with the details provided in Section B of the supplemental material [9]. Our results show that increasing $B$ stabilises the interface, as observed in previous studies [5, 4]. The computations for assessing the influence of $B$ on the interfacial dynamics in the nonlinear regime are then carried out for $F=1.1F_{c}^{N}$ for which square patterns are observed in the surfactant-free case.